The ratio of specific heats is k=Cp/Cv=R, a constant. Combine these two expressions and show that Cv=R/(k-1) and Cp=kR/(k-1)


1
Expert's answer
2021-09-10T00:02:45-0400

We know that


"C_p-C_V=R,\\\space\\
k=\frac{C_p}{C_V}."

Therefore:


"C_V=C_p-R,\\
C_p=kC_V.\\
C_V=kC_V-R,\\
C_V(1-k)=-R,\\\space\\
C_V=\frac{R}{k-1}."

For "C_p", we have


"C_p=R+C_V,\\\space\\
C_V=\frac{C_p}{k}.\\\space\\
C_p=R+\frac{C_p}{k},\\\space\\
C_p\bigg(1-\frac 1k\bigg)=R,\\\space\\
C_p\bigg(\frac {k-1}{k}\bigg)=R,\\\space\\
C_p=\frac{kR}{k-1}."